Suppose that the base year for CPI is 2014. The cost of the CPI basket in 2014 is $50; the cost of the basket in 2015 was $60, and the cost of the basket in 2016 was $66. What was the CPI in 2015?
Answer #1

Answer

CPI=(current year cost of the basket/base year cost of the basket)*100

CPI in 2015=(60/50)*100=120

CPI in 2016=(66/50)*100=132

CPI in 2014=100 as it is the base year